OpenStudy (anonymous):

What is the domain of the expression √8x+16?

OpenStudy (anonymous):

\[What is the domain of the expression \sqrt{8x+16}?\]

OpenStudy (theeric):

I'll make this simple, actually. Just make sure there is no negative under the radical, or the value will be too weird! You want to say that \(8x+16\) is not negative, so \(8x+16\ge 0\). Then solve for \(x\)!

OpenStudy (theeric):

Then you will see all the values that \(x\) can be to make the expression a real number. That's what these questions are asking for! Is that okay?

OpenStudy (anonymous):

Yeah, so 8x+16=0?

OpenStudy (theeric):

It could! There is an acceptable \(x\) value that would let \(8x+16=0\), but there are also many more acceptable values that \(x\) could be. All the values that \(x\) could b is called the "domain."

OpenStudy (theeric):

Have you ever solved inequalities?

OpenStudy (anonymous):

Yea I have

OpenStudy (theeric):

It's pretty much the same as an equation, but, as you might make sense of, multiplying or dividing by a negative number will reverse the direction of the inequality.

OpenStudy (theeric):

I see! So, with \(8x+16\ge0\), you can continue?

OpenStudy (anonymous):

I got x=-2, but how would you get a >, <, or inequality symbols when you already have an equal sign?

OpenStudy (theeric):

Well, you start out with the \(\ge\) sign! That is the whole point, that your expression under the radical (square root thing) has to be \(\ge0\), because it \(\sf can't\) be negative. \(8x+16\ge0\\8x\ge-16\\x\ge-2\) You know?

OpenStudy (anonymous):

Oh, ge is greater than or equal to?

OpenStudy (theeric):

Oh.. On my computer it appears as There's a feature on OpenStudy that might not be working on your computer, then.. I would try to refresh your browser, or try another one! The math feature is very helpful!! Sorry, I didn't know that you couldn't see it!

OpenStudy (theeric):

`\(\ge\)` is the text that is used to generate the \(\ge\) symbol through the website. Are you seeing the same thing twice? Or maybe `'s around the first text?

OpenStudy (anonymous):

YEah same thing twice, but your picture helped thank you and sorry bout that

OpenStudy (theeric):

Don't be, I'm sorry! I hope the math starts working, it's very cool! It does subscripts, exponents, the square root sign over the expression, fractions... I find that Mozilla Firefox works well. Look for a "Math Processing" box at the bottom right. Is it there?

OpenStudy (anonymous):

the yellow box with a pi sign on it?

OpenStudy (theeric):

Hmmm... I don't know what that is.. Maybe the pi near "Mathematics"... I don't know what I can do to help, but I hope the math works for you soon! Are you using a mobile device? That might be why.

OpenStudy (theeric):

I have to go, take care!

OpenStudy (anonymous):

ok thanks I understand the problem now